The susceptibility of magnesium at 300 K is 1.2 × 10^-5. At what temperature will the susceptibility increase to 1.8 × 10^-5?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. x C/T = ⇒ x base 1 / x base 2 = T base / T base 1 ⇒ 1.2 x 10^-5 / 1.8 x 10^-5 = T base 2 / 300 ⇒ T base 2 = 12/18 x 300 = 200 K.